OpenStudy (anonymous):

Which percent is equivalent to 2/5 ? A. 2% B. 5% C. 10% D. 40%

OpenStudy (kidrah69):

if you divide it what do you get?

OpenStudy (kidrah69):

\[2\div5\]

OpenStudy (asad786):

d

OpenStudy (asad786):

srry

OpenStudy (anonymous):

D 20*5=100 20*2=40

OpenStudy (asad786):

free medial

OpenStudy (kidrah69):

medial?

OpenStudy (asad786):

medal

OpenStudy (asad786):

i gave it to @kidrah69

OpenStudy (anonymous):

2/5 = 0.4 and 0.4 = 40%